# Human-AI Collaborative Multi-modal Multi-rater Learning for Endometriosis Diagnosis

### v1

@@ 2.2 Multi-modal Learning
Multi-modal learning has become increasingly crucial in various fields, including medical image analysis and computer vision. It combines data from different sources to provide a more comprehensive understanding of tasks. In medical image analysis, several innovative methods have been developed. These include a chilopod-shaped architecture using modality-dependent feature normalization and knowledge distillation [35], a pixel-wise coherence approach modeling aleatoric uncertainty [36], a trusted multi-view classifier using the Dirichlet distribution [37], and an uncertainty-aware model based on cross-modal random network prediction [38]. Computer vision has seen advancements in multi-modal learning as well. Researchers have combined channel exchanging with multi-modal learning [39], applied self-supervised learning to improve performance [40, 41], enhanced video-and-sound source localization [42], introduced a model for multi-view learning [43], and explored feature disentanglement methods [44, 45].
@@ 2.4 Imaging-based Endometriosis Detection
One crucial indicator to detect endometriosis is the obliteration of the Pouch of Douglas (POD) [7, 8]. However, the development of an AI model that can classify such indicator hinges on the availability of precise POD obliteration annotations from T1/T2 MRIs, a task that is challenging because even experienced clinicians often face uncertainty in identifying this sign. Despite these challenges, there have been some efforts to train multi-modal MRI AI models for POD obliteration classification. For example, Zhang et al. [11] proposed a method to transfer knowledge from ultrasound to MRI for classifying POD obliteration, and Butler et al. [8] explored self-supervised pre-training for multi-modal POD obliteration classification. However, these methods have not been validated against ground-truth labels obtained from surgical reports, making it difficult to determine if their reported accuracy is truly competitive with manual classification.
@@ 4.5 Human-AI Collaborative Multi-modal Multi-rater Ablation Study
5 Qualitative Analyses
We also conduct qualitative analyses about HAICOMM. In Figure 2, (a) and (b) are the input T1 and T2 MRIs, respectively. The table below shows the predictions by the three raters (Rater #1,#2,#3), then the predictions by SSR and ProMix trained with Rater #1’s labels and CROWDLAB’s labels (SSR w GT1, ProMix w GT1, SSR w CL GT, ProMix w CL GT). Next, we show SSR and ProMix trained with CROWDLAB’s labels and relying on human-AI collaborative classification (SSR w HAIC, ProMix w HAIC), followed by the result from our HAICOMM, and the ground truth label from surgical data. The case shows the proposed HAICOMM model can generate correct labels while other models cannot. For 3, the case also shows that the proposed HAICOMM model can generate a correct label while most other methods cannot (only ProMix w GT1 and HAICOMM predict the surgical ground truth label correctly).
Figure 2: Qualitative results of HAICOMM.
Figure 3: Another qualitative result of HAICOMM.

### v2

@@ 2.2 Multi-modal Learning
Multi-modal learning has become increasingly crucial in various fields, including medical image analysis and computer vision. It combines data from different sources to provide a more comprehensive understanding of tasks. In medical image analysis, several innovative methods have been developed. These include a chilopod-shaped architecture using modality-dependent feature normalization and knowledge distillation [35], a pixel-wise coherence approach modeling aleatoric uncertainty [36], a trusted multi-view classifier using the Dirichlet distribution [37], and an uncertainty-aware model based on cross-modal random network prediction [38]. Wang et al. [39, 40, 41] also tried to approach the missing modality issues in the multi-modal learning scenario. Computer vision has seen advancements in multi-modal learning as well. Researchers have combined channel exchanging with multi-modal learning [42], applied self-supervised learning to improve performance [43, 44], enhanced video-and-sound source localization [45], introduced a model for multi-view learning [46], and explored feature disentanglement methods [47, 48].
@@ 2.4 Imaging-based Endometriosis Detection
One crucial indicator to detect endometriosis is the obliteration of the Pouch of Douglas (POD) [7, 8]. However, the development of an AI model that can classify such indicator hinges on the availability of precise POD obliteration annotations from T1/T2 MRIs, a task that is challenging because even experienced clinicians often face uncertainty in identifying this sign [56, 57, 58, 59]. Despite these challenges, there have been some efforts to train multi-modal MRI AI models for POD obliteration classification. For example, Zhang et al. [11] proposed a method to transfer knowledge from ultrasound to MRI for classifying POD obliteration, and Butler et al. [8] explored self-supervised pre-training for multi-modal POD obliteration classification. However, these methods have not been validated against ground-truth labels obtained from surgical reports, making it difficult to determine if their reported accuracy is truly competitive with manual classification.
@@ 4.5 Human-AI Collaborative Multi-modal Multi-rater Ablation Study
4.6 Qualitative Analyses
We also conduct qualitative analyses about HAICOMM. In Figure 2, (a) and (b) are the input T1 and T2 MRIs, respectively. The table below shows the predictions by the three raters (Rater #1,#2,#3), then the predictions by SSR and ProMix trained with Rater #1’s labels and CROWDLAB’s labels (SSR w GT1, ProMix w GT1, SSR w CL GT, ProMix w CL GT). Next, we show SSR and ProMix trained with CROWDLAB’s labels and relying on human-AI collaborative classification (SSR w HAIC, ProMix w HAIC), followed by the result from our HAICOMM, and the ground truth label from surgical data. The case shows the proposed HAICOMM model can generate correct labels while other models cannot. For Figure 3, the case also shows that the proposed HAICOMM model can generate a correct label while most other methods cannot (only ProMix w GT1 and HAICOMM predict the surgical ground truth label correctly).
Figure 2: Qualitative example analysis of HAICOMM.
Figure 3: Another qualitative example analysis of HAICOMM.